When the current has this value, what is the charge on the capacitor?

(6 pts) 4. In the circuit show in the sketch the battery emfis E 48.0 V, R 8.0 n, and C 4.0x10 F. nitially the switch is open and the charge on the capacitor is zero. Attime afler the switch is closed, the current in the resistor is 2.00 A. When the current has this value, what is the charge q on the capacitor? (a) zero (b) 64 10 c 28x10 (d) 192 10 C (e) none of the above answers
Answer #1

Here,

when the curent in the circuit is 2 A

voltage across capacitor , V = E - I * R

V = 48 - 8 * 2

V = 32 V

charge on the capacitor = C * V

charge on the capacitor = 4 *10^-6 * 32

charge on the capacitor = 128 *10^-6 F